# Prethermalization in Fermi-Pasta-Ulam-Tsingou chains

### v1

@@ III.2 First ascending sequence
The second cycle for low-mode excitations starts with two excited modes, $Q_{r}$ and $Q_{2r}$. Evidently, $Q_{r}$ will re-excite $Q_{2r}$, which on the other hand will hit $Q_{4r}$ due to (III). According to (III) mode $Q_{3r=(2r+r)}$ will also be excited. Thus, the second excitation cycle results in the sequence $\{(Q_{r}),(Q_{2r}),(Q_{3r},Q_{4r})\}$. For the third cycle, $Q_{6r}$ and $Q_{8r}$ will be present due to (III), together with all terms that can be formed from $j+k$ sums as dictated by (III). In this ascending phase, terms created from $|j-k|$ are not important, since any $j$ and $k$ will be multiples of $r$ that were already excited in previous cycles. Thus, by induction, the first ascending sequence starting from a root $Q_{r}$, with $2(r-1)<N$, is simply
$\{(Q_{r}),(Q_{2r}),(Q_{3r},Q_{4r}),(Q_{5r},\dots,Q_{8r})$
$\qquad(Q_{9r},\dots,Q_{16r}),(Q_{17r},\dots,Q_{32r}),\dots\}\,$ (14)
until a reflection takes place in (III), i.e. until $j+k=2(N+1)$ for some $x_{j}$ and $Q_{k}$. The reflection, however, will only excite new modes if $2(N+1)-(j+k)$ is not a multiple of $r$, since otherwise the terms would have been already excited during the ascending sequence. This will be explored in the following.
@@ III.3 First descending sequence
    If $r$ divides $2(N+1)$ the first ascending sequence will be reflected into itself (for low modes) or be equivalent to a low mode excitation sequence that reflects into itself (for high modes). This shows, in particular, that $Q_{2}$ is nonthermal for all $N$, since $2(N+1)$ is a always divisible by 2. It also shows that $Q_{1}$ is a trivial thermal root, since it excites all modes in the first ascending sequence for any $N$;
@@
    The reflection takes place when a multiple of $r$ is larger than $N$ for the first time, i.e. when $x_{j}=Q_{(a+1)r}$. Evidently this mode lies outside mode space and the excited mode is actually $Q_{2(N+1)-(a+1)\,r}$ after the reflection. A consequence is that even roots can never excite odd modes, since for even modes $2(N+1)-(a+1)\,r$ is always even. This shows that $Q_{2r}$ are nonthermal roots for all $r$ and all $N$.
@@
The descent continues in multiples of $r$ until the next mode would exit the chain, i.e. until $02(N+1)-(a+b)\,r<$ 0.
@@ III.4 Full excitation sequences
Once the first descending sequence is over, another reflection takes place, now going from left to right, mediated by $\delta_{i,|j-k|}$ in (III). Substituting the last mode number from the descending sequence in this delta, we see that the excited mode after this reflection is $Q_{2(N+1)-(a+b)\,r+r}$. The second ascending sequence will start from this mode number, once again ascending in steps of $r$ until $2(N+1)-(a+b-c)r>N$. Then, we have another reflection from left to right that excites mode $2(N+1)-[2(N+1)-(a+b-c)r-r]=(a+b-c+1)r$, etc. Instead of writing full abstract sequences, we now provide a few examples showing how easy such sequences are to write in practice.
@@
Figure 1: Evolution of mode energies for the root $Q_{3}$ in an $\alpha$-FPUT chain with (a) $N{=}11$ and (b) $N{=}12$ masses. This root is only thermal in the latter case, while in the former its bush is composed of all multiples of $3$ that are smaller than $11$. The mode excitation sequence in (b) is the same as the one predicted using the coupling coefficients, (III.4).
@@
Let us start with a chain with $N=11$ masses. Since $242(N+1)=$ 24, the only three nontrivial thermal roots in this chain are $Q_{5}$, $Q_{7}$ and $Q_{11}$. The root $Q_{3}$ is evidently nonthermal since it divides 24, but let’s take a look at its excitation sequence. The first ascending sequence is $\{(Q_{3}),(Q_{6}),(Q_{9})\}$, and we stop here because the next term falls outside mode space. The reflection term comes from $i=2(N+1)-(6+9)=9$, which is already excited. This will be now propagated backwards to $Q_{6}$ and $Q_{3}$ accoding to (III), retracing the ascending sequence and never exciting any new mode. Now, for the leftmost reflection, the excited mode is $Q_{|3-3|}$, which does not exist. Thus, the bush of $Q_{3}$ is given by $\{(Q_{3}),(Q_{6}),(Q_{9})\}$. This is verified numerically in Fig. 1(a).
Figure 2: Evolution of mode energies for the first five odd modes, namely: (a) $Q_{1}$; (b) $Q_{3}$; (c) $Q_{5}$; (d) $Q_{7}$; (e) $Q_{9}$; (d) and a random initial condition, all with energy density $h{\approx}0.085$ in an $\alpha$-FPUT chain with $N{=}63$. Line colors correspond to mode frequencies, going from lower (blue) to higher (green), with their equilibrium value $h{=}E/N$ shown as a black dashed line. Equipartition times (orange triangles) are defined as the instant the time-evolving spectral entropy, shown in black in the inset, first attains its equilibrium value $\langle\eta\rangle$, given by the horizontal red line.
@@
Let us now consider what happens to $Q_{3}$ for $N=12$ instead. The first ascending sequence is now updated to $\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12})\}$. The reflection gives $i=2(N+1)-(3+12)=11$, which was not previously excited and is not a multiple of $r=3$. This will then propagate backwards to $11-3=8$, $11-6=5$, etc. The full excitation sequence is then given by
$\nearrow$ $\quad\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12},Q_{11=26-(12+3)}),$
[…]
$\nearrow$ $\quad(Q_{4},Q_{7},Q_{10})\}\,,$ (16)
[…]
As another example, for $Q_{3}$ and $N=13$ we have
$\nearrow$ $\quad\{(Q_{3}),(Q_{6}),(Q_{9},Q_{12},Q_{13=28-(12+3)}),$
[…]
$\nearrow$ $\quad(Q_{2=|1-3|},Q_{5},Q_{8},Q_{11})\}\,.$ (17)
In both (III.4) and (III.4) the root $Q_{3}$ resulted in three excitation sequences: two ascending and one descending. If $Q_{5}$ is chosen as the root, then the full excitation sequence for $N=13$ is
$\nearrow$ $\quad\{(Q_{5}),(Q_{10},Q_{13=28-(10+5)}),$
[…]
$\nearrow$ $\quad(Q_{7},Q_{12},Q_{11=28-(12+5)}),$
[…]
$\nearrow$ $\quad(Q_{4=|1-5|},Q_{9})\}\,.$ (18)
In general, it is easy to see that since every step in the sequence is performed in multiples of $r$, one needs $r$ ascents and descents (or $r-1$ reflections) to cover mode space for a thermal root $Q_{r}$, no matter what $N$ is chosen. Since large $r$ requires more cycles to cover mode space, it means less modes are excited at the end of each descending or ascending sequence. One can once again use the case of $N=13$ as an example: Here, at the end of the first ascending sequence, the root $Q_{1}$ has already excited all modes in the chain, which is a unique property of the fundamental root; $Q_{3}$, however, has only excited 5 of the 13 available modes, and $Q_{5}$ only 3 out of 13. In Sec. IV we will show that this effect strongly impacts any quantity computed from evolving roots.
@@ IV Numerical experiments
Figure 3: Mode energies, $E_{j}$, as a function of mode number, $j$, for roots $Q_{1}$, $Q_{3}$ and $Q_{9}$ for fixed times, with their equilibrium (average) value, $h$, shown as a black dashed line. Each panel is a vertical slice of a panel in Fig. 2 at (a) $t{=}\tau_{\mathrm{eq}}/100$; (b) $t{=}\tau_{\mathrm{eq}}$ and (c) $t{=}10\tau_{\mathrm{eq}}$ (note that $\tau_{\mathrm{eq}}$ is different for each root). The short time of (a) allows for a clear visualization of the energy cascade flowing between modes that are multiples of 1, 3 and 9, as described in Sec. III. At equipartition most of the energies per site have the same order of magnitude, as seen in (b), with stronger mixing still taking place at high energy modes while the lower half of the excitation sequence remains slightly isolated from the other modes. Even after ten equipartition times the nearly homogeneous energy distribution hints at which root was initially excited, as can be seen in the above-average energies of modes $Q_{9}$ and $Q_{18}$ for the root $Q_{9}$ in (c).
In this section we perform computations in the $\alpha$-FPUT chain (1) with $N{=}63$ sites. This value of $N$ is chosen because $1282(N+1)=$ 128 is not divisible by any odd integer other than 1, such that all odd roots are thermal for this chain. Although initial states, final propagation times and energy densities will vary, evolution is always computed using an optimized second order symplectic integrator [37] with step size $\Delta t=0.2$. This results in a maximum relative energy error of around $10^{-3}$ for the largest energy densities chosen, and much smaller for most of them. We will focus on the evolution of random initial conditions and the first five odd roots $Q_{1}$, $Q_{3}$, $Q_{5}$, $Q_{7}$ and $Q_{9}$.
@@ IV.1 Metastable plateaus
Sec. III described how energy can remain localized within a small subset of modes before spreading throughout mode space: The higher excited the root, the smaller the initial subset of excited modes, and also the slower the energy transfer. During this transfer time the dynamics is mostly localized within the excited modes, which essentially “wait” while slowly transferring energy to the ones in the next cycles, forming the metastable plateaus well known in FPUT literature [17, 38]. The longer it takes for the energy to reach the end of the sequence, the longer a metastable plateau survives.
@@
The well-known fact that the metastable plateaus associated to roots become more visible at lower energy densities is unsurprising, since in the limit of zero energy density such modes are normal modes of the associated harmonic lattice. However, the behavior of plateaus for excited roots was not previously addressed. To this end, Fig. 2 shows a comparison of evolving mode energies and spectral entropies for the prototypical $\alpha$-FPUT chain (1) computed starting from the first 5 odd roots in a chain with $N=63$ masses, together with a random initial state. Fig. 2(a) displays the standard mode energy dynamics found in literature: An excitation of $Q_{1}$ monotonically spreads to all other modes and eventually results in energy equipartition, with all $E_{k}$’s converging to the mean energy $E/N=h$ (black dashed line). Metastable plateaus are clearly visible in the prethermal regime, in which a measurement would indicate that the system is not ergodic and cannot be described by equilibrium statistical mechanics. However, such a statement is evidently incorrect, as waiting a bit longer would result in a system that does display statistical behavior.
@@
The approach from an out-of-equilibrium initial state towards a thermal one is most clearly seen in tracking the spectral entropy as a function of time, shown in the insets: The moment it touches the red line, which represents its equilibrium value (8), is where we consider equipartition to have been achieved. The associated equipartition times are displayed in the main plots as an orange triangle and slowly increase as we move from $Q_{1}$ up to $Q_{9}$, with the shortest time corresponding to the random initial condition. Evidently, this latter case is less interesting since the equipartition time depends strongly on the initial sampling, but for roots the structures formed by mode energies as they evolve in time show that, indeed, higher-excited roots isolate an increasingly smaller subset of modes in mode space. Fig. 2(a) also shows a dense subset of modes homogeneously concentrated around the mean, while Fig. 2(e) clearly shows that two modes, namely $Q_{9}$ and $Q_{18}$, remain relatively isolated from the others for much longer times. To see this more clearly, in Fig. 3 we take vertical slices of Fig. 2 at three different fixed times for $Q_{1}$, $Q_{3}$ and $Q_{9}$. At $t{=}\tau_{\mathrm{eq}}$ low frequency modes that are multiples of 3 and 9 can still be seen to have above-average energies for roots $Q_{3}$ and $Q_{9}$. At $t{=}10\tau_{\mathrm{eq}}$ this effect has already dissipated for $Q_{3}$, but the energies of modes $Q_{9}$ and $Q_{18}$ remain slightly above the mean energy when evolving $Q_{9}$, as visible in Fig. 3(c).
@@
Evidently, the longer it takes for a subset of modes to reach equipartition, the longer the prethermal regime. This is clearly seen in Fig. 2: The relatively obfuscated metastability associated to the $Q_{1}$-root seen in panel (a) is already maximized in (b), where $Q_{3}$ is evolved, and becomes more and more blatant as higher-excited modes are used as roots. Evidently, the random initial condition in Fig. 2(e) does not display metastability, since here all modes are excited at once and the slow, selection-rule mediated transferring of energy from excited modes to previously unexcited ones does not take place. Since the mean energy is a constant of motion, the prethermal regime’s duration can also be visualized by plotting the standard deviation of $E_{k}(t)$, $\sigma[E_{k}(t)]$, as a function of time, which will be approximately constant while the energy is still strongly localized in a subset of normal modes. This is indeed seen to be true in Fig. 4, where it is also clear that the metastability’s lifetime increases monotonically as a function of root excitation for this energy density.

### v2

@@ III.2 First ascending sequence
The second excited mode is obtained from $Q_{r}$ and $Q_{2r}$. Evidently, $Q_{2r}$ excites $Q_{4r}$ due to (III) and, since $Q_{2r}$ is of $\mathcal{O}(\alpha^{1})$, the amplitude of $Q_{4r}(t)$ goes as $\alpha\,Q_{2r}Q_{2r}=\mathcal{O}(\alpha^{4})$ for short times. Mode $Q_{3r{=}r+2r}$ is also excited according to (III), but now the short-time amplitude will scale as $\alpha\,Q_{r}Q_{2r}=\mathcal{O}(\alpha^{2})$. By induction, the first ascending sequence for a root $Q_{r}$ is given by
$\{Q_{r},Q_{2r},Q_{3r},Q_{4r},Q_{5r},\dots,Q_{ar}\}\,.$ (14)
It includes all multiples of the root mode ordered by how their short-time dynamics scales with $\alpha$, namely $\mathcal{O}(\alpha^{a-1})$, until a reflection takes place in (III), i.e., until $j{+}k{=}2(N{+}1)$ for some $Q_{j}$ and $Q_{k}$. The reflection, however, will only excite new modes if $2(N{+}1){-}(j{+}k)$ is not a multiple of $r$, since otherwise the terms would have already been excited during the ascending sequence.
@@ III.3 First descending sequence
    If $r$ divides $2(N{+}1)$, then the first ascending sequence is reflected into itself. This shows, in particular, that $Q_{2}$ is nonthermal for all $N$, since $2(N{+}1)$ is a always divisible by 2 [39, 40]. It also shows that $Q_{1}$ is a trivial thermal root, since it excites all modes in the first ascending sequence for any $N$;
@@
    The reflection takes place when a multiple of $r$ is larger than $N$ for the first time, i.e., when $Q_{j}=Q_{(a+1)r}$. Evidently this mode lies outside mode space and the excited mode is actually $Q_{2(N+1)-(a+1)\,r}$ after the reflection. A consequence is that even roots can never excite odd modes, since for even modes $2(N{+}1){-}(a{+}1)\,r$ is always even [40]. This shows that $Q_{2r}$ are nonthermal roots for all $r$ and all $N$.
@@
The descent continues in multiples of $r$ until the next mode would exit the chain, i.e., until $02(N+1)-(a+b)\,r<$ 0. Despite the reflection, it is important to note that each term in the descending sequence comes from a product like $Q_{ar}Q_{br}$, and will therefore have an amplitude that scales perturbatively as $\mathcal{O}(\alpha^{a+b-1})$ for short times, just as in the ascending sequence. This shows both ascending and descending sequences can be tied together and monotonically ordered based on how their short-time solutions scale with $\alpha$.
@@ III.4 Full excitation sequences
Once the first descending sequence is over, another reflection takes place, now going from left to right, mediated by $\delta_{i,|j-k|}$ in (III). Substituting the last mode number from the descending sequence in this delta, we see that the excited mode after this reflection is $Q_{2(N+1)-(a+b)\,r+r}$. The second ascending sequence will start from this mode number, once again ascending in steps of $r$ until $2(N+1)-(a+b-c)r>N$. Then, we have another reflection from left to right that excites mode $2(N+1)-[2(N+1)-(a+b-c)r-r]=(a+b-c+1)r$, and so on. Instead of writing full abstract sequences, we now provide a few examples showing how easy it is to write them in practice.
@@
Figure 1: Evolution of mode energies for the root $Q_{3}$ in an $\alpha$-FPUT chain with (a) $N{=}11$ and (b) $N{=}12$ masses, both for $h{\approx}0.16$ and $\alpha{=}1/4$. The root $Q_{3}$ is only thermal in the latter case, while in the former its bush is composed of all multiples of $3$ that are smaller than $11$. The mode excitation sequence in (b) is the same as the one predicted using the coupling coefficients, (III.4).
@@
Let us start with a chain with $N{=}11$ masses. Since $242(N{+}1){=}$ 24, the only three nontrivial thermal roots in this chain are $Q_{5}$, $Q_{7}$ and $Q_{11}$. The root $Q_{3}$ is evidently nonthermal since it divides 24 (c.f. Sec. III.3), but let’s take a look at its excitation sequence. Denoting a sequence starting from root $r$ in a system with $N$ normal modes by $S_{r}^{N}$, the first ascending sequence is $S_{3}^{11}=\{Q_{3},Q_{6},Q_{9}\}$, and we stop here because the next term falls outside mode space. The reflection term comes from $i{=}2(N{+}1){-}(6{+}9)=9$, which is already excited. This will be now propagated backwards to $Q_{6}$ and $Q_{3}$ accoding to (III), retracing the ascending sequence and never exciting any new mode. Now, for the leftmost reflection, the excited mode is $Q_{|3-3|}$, which does not exist. Thus, the bush of $Q_{3}$ is given by $\{Q_{3},Q_{6},Q_{9}\}$. This contradicts a statement made in [40], which states that all odd modes should be excited if $r$ and $N$ are relatively prime, but is verified numerically in Fig. 1(a).
@@
Let us now consider what happens to $Q_{3}$ for $N{=}12$ instead. The first ascending sequence is now updated to $\{Q_{3},Q_{6},Q_{9},Q_{12}\}$. The reflection gives $i{=}2(N{+}1){-}(12{+}3){=}11$, which was not previously excited and is not a multiple of $r$. The next excited modes will then be $11{-}3{=}8$, $11{-}6{=}5$, etc, and the full excitation sequence is
$S_{3}^{12}$ =$=\{Q_{3},Q_{6},Q_{9},Q_{12},Q_{11=26-(12+3)},$
$\qquad Q_{8},Q_{5},Q_{2},Q_{1=|2-3|},Q_{4},Q_{7},Q_{10}\}\,,$ (16)
where modes are ordered by how their short-time solutions scale with $\alpha$. This excitation sequence is seen to match numerics in Fig. 1(b).
As another example, for $S_{3}^{13}$ we have
$S_{3}^{13}$ =$=\{Q_{3},Q_{6},Q_{9},Q_{12},Q_{13=28-(12+3)},$
$\qquad Q_{10},Q_{7},Q_{4},Q_{1},Q_{2=|1-3|},Q_{5},Q_{8},Q_{11}\}\,,$ (17)
and, for $S_{5}^{13}$,
$S_{3}^{13}$ =$=\{Q_{5},Q_{10},Q_{13=28-(10+5)},Q_{8},Q_{3},Q_{2=|3-5|},$
$\qquad Q_{7},Q_{12},Q_{11=28-(12+5)},$
$\qquad Q_{6},Q_{1},Q_{4=|1-5|},Q_{9}\}\,.$ (18)
III.5 Mode connections and amplitudes
Excitation sequences such as (III.4) and (III.4), and also the trivial sequence generated by $Q_{1}$, completely fill mode space for $N{=}13$. However, these sequences do not provide insight into how energy is exchanged between excited modes – an aspect we will briefly explore in the following.
Starting from (III.4), we see that the pair that most strongly connects to $Q_{13}$ in $S_{5}^{13}$ is $(Q_{5},Q_{10})$, since the right-hand side of (11) gives $\alpha\,Q_{5}Q_{10}=\mathcal{O}(\alpha^{2})$. There are also connections formed by other pairs, such as $Q_{5}$ and $Q_{8}$, but these connections are of higher order in $\alpha$, e.g., $\alpha\,Q_{8}Q_{5}=\mathcal{O}(\alpha^{4})$. Thus, the short-time regime will see $Q_{13}$ exchanging energy primarily with $Q_{5}$ and $Q_{10}$. The situation is similar in $S_{3}^{13}$ in (III.4), where $(Q_{12},Q_{3})$ and $(Q_{9},Q_{6})$ both connect to $Q_{13}$ and have weighs of $\mathcal{O}(\alpha^{4})$. The case of $S_{1}^{13}$ is rather extreme, since the sum or difference of any two mode numbers that equals 13 or 15 all connect to $Q_{13}$, and all connections scale as $\mathcal{O}(\alpha^{12})$.
The argument above sheds light on why energy remains localized in a subset of modes for short times, but the simplified form of (11) glosses over the multiplier $\omega_{i}\omega_{j}\omega_{k}$ in (II.1). These frequencies $\omega$, as seen in (3), increase monotonically for higher mode numbers. A consequence is that, despite $Q_{3}$ in $S_{1}^{13}$ and $Q_{13}$ in $S_{5}^{13}$ both scaling perturbatively as $\mathcal{O}(\alpha^{2})$, the amplitude of $Q_{13}$ in $S_{5}^{13}$ will lie above the one of $Q_{3}$ in $S_{1}^{13}$ for short times if the roots are excited with the same energy density. The same conclusion could be drawn from noticing that, for a fixed energy density, energy has to be shared among a smaller subset of excited modes in $S_{5}^{13}$ when compared to $S_{1}^{13}$ in the short-time regime, rendering local energies larger. These larger energies take longer to spread to other modes, since $Q_{13}$ in $S_{5}^{13}$ has exactly the same number of connections as $Q_{3}$ in $S_{1}^{13}$ (and they are also of the same order). Note also that the next thermal root in $S_{r}^{13}$ would be $r{=}7$, but this mode is a q-breather for $N{=}13$. Thus, increasing mode number is tantamount to approaching a state for which energy does not spread in mode space, which offers a more qualitative way of interpreting the slowing down of energy equipartition for increasing $r$.
@@ IV Numerical experiments
Figure 2: Evolution of mode energies for the first five odd modes, namely: (a) $Q_{1}$; (b) $Q_{3}$; (c) $Q_{5}$; (d) $Q_{7}$; (e) $Q_{9}$; and (f) a random initial condition, all with energy density $h{=}H(\boldsymbol{p},\boldsymbol{x})/N{\approx}0.085$
in an $\alpha$-FPUT chain with $N{=}63$ and $\alpha{=}1/4$. Line colors correspond to mode frequencies, going from lower (blue) to higher (green), with their equilibrium value shown as a black dashed line. Equipartition times (orange triangles) are defined as the instant the time-evolving spectral entropy, shown in black in the inset, first attains its equilibrium value $\langle\eta\rangle$, given by the horizontal red line.
Figure 3: Mode energies, $E_{j}$, as a function of mode number, $j$, for roots $Q_{1}$, $Q_{3}$ and $Q_{9}$ for fixed times, with their equilibrium (average) value, $h$, shown as a black dashed line. Each panel is a vertical slice of a panel in Fig. 2 at (a) $t{=}\tau_{\mathrm{eq}}/100$; (b) $t{=}\tau_{\mathrm{eq}}$ and (c) $t{=}10\tau_{\mathrm{eq}}$ (note that $\tau_{\mathrm{eq}}$ is different for each root). The short time of (a) allows for a clear visualization of the energy cascade flowing between modes that are multiples of 1, 3 and 9, as described in Sec. III. At equipartition most of the site energies have the same order of magnitude, as seen in (b), with stronger mixing still taking place at high mode numbers while the lower half of the excitation sequence remains slightly isolated from the other modes. Even after ten equipartition times the nearly homogeneous energy distribution hints at which root was initially excited, as can be seen in the above-average energies of modes $Q_{9}$ and $Q_{18}$ for the root $Q_{9}$ in (c).
In this section we perform computations in the $\alpha$-FPUT chain (1) with $N{=}63$ sites. This value of $N$ is chosen because $1282(N{+}1){=}$ 128 is not divisible by any odd integer other than 1, such that all odd roots are thermal for this chain. Although initial states, final propagation times and energy densities will vary, evolution is always computed using an optimized second order symplectic integrator with step size $\Delta t{=}0.2$ [42]. This results in a maximum relative energy error of around $10^{-3}$ for the largest energy densities chosen, and much smaller for most of them. We will focus on the evolution of random initial conditions and the first five odd roots $Q_{1}$, $Q_{3}$, $Q_{5}$, $Q_{7}$ and $Q_{9}$. The sampling of all initial conditions starts by setting $\boldsymbol{x}(0){=}0$. Then, for random states, each momentum component is drawn from a Maxwell-Boltzmann distribution, while for mode $Q_{k}$ we have momentum components fixed as $p_{j}(0){=}\sin[\pi kj/(N{+}1)]$. Once sampled, momenta are then renormalized in order to achieve the desired energy density for both types of initial conditions.
@@ IV.1 Metastable plateaus
Sec. III described how energy can remain localized within a small subset of modes before spreading throughout mode space: the higher excited the root, the smaller the subset of excited modes and the larger their amplitudes, such that more time is needed in order to distribute their energy among other modes. During this transfer time the dynamics is mostly localized within the subset of excited modes, which essentially “wait” and form the metastable plateaus prominent in FPUT literature [18, 43, 14].
@@
The well-known fact that the metastable plateaus associated to roots become more visible at lower energy densities is unsurprising, since in the limit of zero energy density such modes are normal modes of the associated harmonic lattice. However, the behavior of plateaus for excited roots was not previously addressed. To this end, Fig. 2 shows a comparison of evolving mode energies and spectral entropies for the $\alpha$-FPUT chain and initial states described in the introduction to this section. Fig. 2(a) displays the standard mode energy dynamics found in literature: energy initially placed in $Q_{1}$ monotonically spreads to all other modes and eventually results in energy equipartition, with all $E_{k}$’s converging to the mean energy $h=H(\boldsymbol{p},\boldsymbol{x})/N$ (black dashed line). Metastable plateaus are clearly visible in the prethermal regime, in which a measurement would indicate that the system is not ergodic and cannot be described by equilibrium statistical mechanics. This statement is evidently incorrect, as waiting longer would result in a system that does display statistical behavior.
@@
The approach from an out-of-equilibrium initial state towards a thermal one is most clearly seen in tracking the spectral entropy as a function of time, shown in the insets: the moment it touches the red line, which represents its equilibrium value (8), is where we consider equipartition to have been achieved. The associated equipartition times are displayed in the main plots as an orange triangle and slowly increase as we move from $Q_{1}$ up to $Q_{9}$, with the shortest time corresponding to the random initial condition. Evidently, this latter case is less interesting since the equipartition time depends strongly on the initial sampling, but for roots the structures formed by mode energies as they evolve in time show that, indeed, higher-excited roots isolate an increasingly smaller subset of modes in mode space. Fig. 2(a) also shows a dense subset of modes homogeneously concentrated around the mean, while Fig. 2(e) clearly shows that two modes, namely $Q_{9}$ and $Q_{18}$, remain relatively isolated from the others for much longer times. To see this more clearly, in Fig. 3 we take vertical slices of Fig. 2 at three different fixed times for $Q_{1}$, $Q_{3}$ and $Q_{9}$. At $t{=}\tau_{\mathrm{eq}}/100$ the contents of Sec. III.5 can be clearly visualized, namely the fact that increasing root number traps energy in a smaller subset of modes with higher amplitudes in the short-time regime. At $t{=}\tau_{\mathrm{eq}}$ low frequency modes that are multiples of 3 and 9 can still be seen to have above-average energies for roots $Q_{3}$ and $Q_{9}$. At $t{=}10\tau_{\mathrm{eq}}$ this effect has already dissipated for $Q_{3}$, but the energies of modes $Q_{9}$ and $Q_{18}$ remain slightly above the mean energy when evolving $Q_{9}$, as visible in Fig. 3(c).
@@
Evidently, the longer it takes for a subset of modes to reach equipartition, the longer the prethermal regime. This is clearly seen in Fig. 2: The relatively obfuscated metastability associated to the $Q_{1}$-root seen in panel (a) is already maximized in (b), where $Q_{3}$ is evolved, and becomes more and more blatant as higher-excited modes are used as roots. Evidently, the random initial condition in Fig. 2(f) does not display metastability, since here all modes are excited at once and the slow, selection-rule mediated transferring of energy from excited modes to previously unexcited ones does not take place. Since the mean energy is a constant of motion, the prethermal regime’s duration can also be visualized by plotting the time-dependent standard deviation of mode energies in (2),
$\sigma_{E}(t)=\sqrt{\frac{1}{N-1}\sum_{j=1}^{N}\left(E_{j}(t)-h\right)^{2}}\,,$ (19)
which will be approximately constant while the energy is still strongly localized in a subset of normal modes. This is indeed seen to be true in Fig. 4, where it is also clear that the metastability’s lifetime increases monotonically as a function of root number for this energy density.